### v3

# AM-RADIO: Agglomerative Model – Reduce All Domains Into One If you can’t beat them, join them (and then beat them)

Abstract
A handful of visual foundation models (VFMs) have recently emerged as the backbones for numerous downstream tasks. VFMs like CLIP, DINOv2, SAM are trained with distinct objectives, exhibiting unique characteristics for various downstream tasks. We find that despite their conceptual differences, these models can be effectively merged into a unified model through multi-teacher distillation. We name this approach AM-RADIO (Agglomerative Model – Reduce All Domains Into One). This integrative approach not only surpasses the performance of individual teacher models but also amalgamates their distinctive features, such as zero-shot vision-language comprehension, detailed pixel-level understanding, and open vocabulary segmentation capabilities. In pursuit of the most hardware-efficient backbone, we evaluated numerous architectures in our multi-teacher distillation pipeline using the same training recipe. This led to the development of a novel architecture (E-RADIO) that exceeds the performance of its predecessors and is at least 7x faster than the teacher models. Our comprehensive benchmarking process covers downstream tasks including ImageNet classification, ADE20k semantic segmentation, COCO object detection and LLaVa-1.5 framework.
Code: https://github.com/NVlabs/RADIO.
**footnotetext: Equal contribution
Results
Framework
Figure 1: AM-RADIO - is a multi-teacher distillation framework that efficiently trains new vision foundation models of arbitrary architecture. It unifies unique attributes (like zero-shot text grounding, dense correspondence) of each teacher into a single model that even outperforms them on majority of the tasks.
1 Introduction
Knowledge Distillation [23] has been a very successful and popular technique for transferring the knowledge of a “teacher” model into a typically smaller “student” model. In the original formulation, both the student and the teacher operate on the same in-domain dataset, and the student simultaneously matches the logits of the teacher, and the GT labels. Instead of using labeled images, an alternative approach is to train the student model to match the features of the teacher model [48, 25, 1, 22, 65, 51, 56].
Alternative to using a smaller student model, [58] employ an iterative learning procedure with a high-capacity model where the “student” of equal or greater capacity than the teacher is trained with heavy augmentation applied to the student. Once trained, they expand the dataset by pseudo-labeling new data using the trained student. They then make the student become the teacher, and repeat the process. An important finding in this work is that the student is capable of surpassing the performance of the teacher.
Again in [23], the authors explore the concept of ensemble distillation, where there are multiple teachers, each of which having restricted domain knowledge. [69] provides an overview of multi-teacher distillation, and proposes that instead of matching the summary of an ensemble of teachers, the student can match the features of each individual teacher via some learned non-shared mapping from the representation space of the student to each teacher. Of interest in their approach is that the student and teacher don’t need to share the same architecture, and also that treating teachers individually yields improved performance.
Recently, the concept of Foundation Models (FMs) [3] has emerged, with the general understanding that these models are large, accurate, and typically very expensive to train. They are meant to represent models that, through training on very large datasets, are broadly applicable to numerous downstream tasks. A seminal example of such models is CLIP [46], which trains on web-scale weakly supervised (image, caption) pairs, and results in exceptional zero-shot performances on a wide array of computer vision benchmarks, such as 76.2% zero-shot top-1 accuracy on ImageNet-1K. While CLIP is firmly a FM, another model, DINOv2 [43] has emerged with broad capabilities, often surpassing CLIP. In particular, for tasks that involve holistic representations for an entire image (e.g., classification), DINOv2 isn’t a clear victor compared to CLIP. However, if we turn our attention to dense tasks (e.g., semantic segmentation), DINOv2 shows clear superiority to CLIP on benchmarks such as ADE20k [68] and Pascal VOC [15]. Separately, SAM (Segment Anything) [31] is gaining popularity for its excellent open-vocabulary semantic segmentation abilities, which we hypothesize means it has strong dense feature representations.
We introduce AM-RADIO (Agglomerative Model – Reduce All Domains Into One) with the goal of learning from multiple foundational models simultaneously. We observe that, when given a student model of sufficient capacity, it is often able to exceed any of its teachers on important metrics such as K-NN ImageNet Top-1 accuracy [57, 8, 43] (refer to [57] for implementation details). In addition to performing well on representative foundational benchmarks, by virtue of the training framework, our student models are able to mimic their teacher models, and thus are able to perform downstream tasks that are otherwise performed by the teachers. Examples of this include CLIP-ZeroShot applications, since the language model trained by CLIP is compatible with our student, and also Segment-Anything tasks, as the student is able to replace the vision encoder and interface with the already-trained mask decoders.
In addition to the student itself being a large (and slow) model, we study the effect of using a more hardware-efficient model architecture. Most works are not directly comparable as they use different training recipes, even when evaluated on the same dataset such as ImageNet-1k. To this end, we evaluate more than 10 promising architectures under the same training recipe for a direct comparison. We reveal that CNN-like architectures are faster but less expressive than ViTs. This has led to the development of a novel hybrid architecture, E-RADIO, that exceeds the performance of its predecessors and is at least 7x faster than teacher models.
Our main contributions are as follows:
  • We describe a general methodology for distilling multiple distinct foundation models into one, including models with incompatible input resolutions.
  • We show that these student models are able to outperform their teachers on representative benchmarks.
  • We demonstrate that these student models can either drop-in replace their teachers, or their features can be used directly in downstream applications such as providing visual encoding for LLaVA [36, 35].
  • We benchmark a number of efficient architectures and propose a new architecture (E-RADIO) that allows for similar model quality at significant speedups.
2 Related Work
Knowledge Distillation
The underpinning of our work is based on the method of Knowledge Distillation [23, 30, 4, 42, 5] which aims to train a “student” model using soft targets produced by an already-trained “teacher” model, using the the teacher’s output logits as “soft” labels. Alternatively, distillation can be performed using intermediate network activations [48, 25, 1, 22, 65, 51, 56]. In general, due to the heterogeneous nature of the different teacher foundation models that we employ, we ignore any potential labels coming from the data, and we ignore the logits of teachers, and simply opt to match the feature representations of the teachers before any task-specific processing stages.
Multi-Teacher Distillation
There is also a body of work that studies distilling a student model jointly from multiple teacher models simultaneously [23, 37, 69, 64, 66, 61, 45, 62, 32, 2, 17]. Because of the heterogeneous domains that our teacher models cover, we don’t apply approaches that marginalize teachers into a unified label, and instead map students to each teacher independently using teacher-specific projection heads from the unified student representation. Although the reason behind this method in [69] is different, we find the same overall strategy to be effective. While [56] doesn’t study matching the features of multiple teachers simultaneously, we are able to extend their paradigm via the different projection heads. To preserve drop-in compatibility with teacher frameworks, we eliminate the feature normalization in the loss function.
Distilling Foundation Models
Foundation Models [3] are meant to be generalist models that are trained on massive amounts of data, and are typically resource intensive to train from scratch. In the vein of single-teacher distillation, [43] employ self-distillation to train their smaller variants from the larger teacher. [56] distills their model from a CLIP [46] teacher. Instead of focusing our energy on one teacher in particular, we instead grab high-quality versions of the popular CLIP [46] (using OpenCLIP [27]) model, DINOv2 [43], and SAM [31]. Concurrently with our work, [55] describe a methodology for merging a CLIP model into a pretrained SAM model via distillation, which is, in spirit, quite similar to this approach. In contrast to theirs, we include DINOv2 and also simplify the objective to straight-forward feature matching. Since we don’t rely on the student model to be pre-trained, it also gives us the flexibility to have the student be an architecture distinct from any teacher.
3 Knowledge Agglomeration
We propose a framework to train a vision foundation model from scratch via multi-teacher distillation as shown in Figure 1. We assume that each teacher brings unique properties to the foundational vision model, and the resulting trained model will integrate these attributes.
3.1 Overview
As an initial assumption, we expect that the teacher models are reasonably capable of representing a broad swath of images found on the internet, coming from datasets such as ImageNet (1k or 21k) [13], LAION-400M [49] or DataComp-1B [18]. With this in mind, we choose to study 3 seminal teacher models: CLIP [46], DINOv2 [43], and SAM [31] as they have demonstrated outstanding performance over a broad range of tasks (as in CLIP), or specifically strong performance on downstream dense tasks, such as semantic segmentation under linear probe (as in DINOv2), or open-vocabulary segmentation (as in SAM). Because these teacher models come from such diverse domains, we omit any form of supplemental ground truth guidance and treat the aforementioned datasets simply as sources of images. To assess the quality of our models, we adopt a set of representative metrics across a few broad domains.
  • Image level reasoning: (i) k-NN Top-1 accuracy on ImageNet-1K, and (ii) Zero-Shot accuracy using the CLIP teacher’s language model. ”Zero-Shot” refers to CLIP-style classification using prompt engineering with the text encoder (not available for DINOv2). k-NN [57, 8, 43] embeds the model’s summary feature vector for every image in the training set, and then for each validation image, it uses a weighted sum of the $k$ nearest training vectors to elect a label.
  • Pixel-level visual tasks: segmentation mIOU on (i) ADE20K and (ii) Pascal VOC - under the linear probe setting, details in Section 5.5.
  • Large Vision-Language Models: we plug our frozen vision encoder model into LLaVA [36] and evaluate it on a wide set of tasks including GQA [26], TextVQA [50], ScienceQA [41] and VQAv2 [20]. Details in Section 5.6.
3.2 Distillation Dataset Choice
In table 1 we study the effect of different datasets on downstream metrics. While the highest image classification metrics are achieved using ImageNet-1K as the training dataset, we argue that it doesn’t fairly measure “zero shot” performance as the student directly learns the teacher features in the evaluation domain. For this reason, we opt for the DataComp-1B dataset.
Config Dataset k-NN Zero Shot ADE20K
A Teacher Best 83.41 77.19 47.53
ImageNet 1K 84.79 77.93 48.00
ImageNet 21K 84.03 77.25 48.70
LAION-400M 82.84 74.85 48.41
DataComp-1B 83.17 76.90 48.42
B Teacher Best 83.41 80.51 47.53
ImageNet 1K 84.79 80.44 48.11
ImageNet 21K 84.61 80.10 48.65
LAION-400M 83.77 77.46 48.6
DataComp-1B 83.91 78.51 49.01
Table 1: Ablation study on the choice of different training datasets. Configuration A: we use DINOv2 ViT-G/14 and OpenCLIP ViT-H/14 teachers and a ViT-L/14 [14] student model with CPE [29]; B: we replace the CLIP teacher with MetaCLIP ViT-H/14 and the student model with ViT-H/14. Both “k-NN” and “Zero Shot” are for ImageNet-1k. ADE20k refers to mIOU linear probe on ADE20k. We also report the maximum metrics coming from either of the teacher models.
3.3 Loss Formulation
Because we don’t have (broadly applicable) ground truth data for each teacher for each image, we instead opt to match the features coming from each teacher’s vision encoder. In particular, we distinguish between the summary feature vector and the spatial feature vectors for each teacher. Because all of our teachers use the ViT architecture, the spatial features are the feature vectors for each input patch coming from the final layer of the model. The summary feature is computed differently based on the model. For CLIP (we use OpenCLIP [27]) and DINOv2, we use the “class token” as the summary feature vector, and for SAM we use the average over the spatial features.
Let $f\left(x|\Theta_{0}\right)$ be the student vision encoder with parameters $\Theta_{0}$, and $y_{i}^{s}=h_{i}^{(s)}(x^{\prime}|\Theta_{i}^{(s)})$ be the learned student head matching teacher summary features $z_{i}^{(s)}=t_{i}^{(s)}\left(x|\Phi_{i}\right)$ with student adaptor parameters $\Theta_{i}^{(s)}$ and teacher parameters $\Phi_{i}$.
$x^{\prime}$ $=f\left(x|\Theta_{0}\right);$ $y_{i}^{(s)}$ $=h_{i}^{(s)}\left(x^{\prime}|\Theta_{i}^{(s)}\right);$ (1)
$z_{i}^{(s)}$ $=t_{i}^{(s)}\left(x|\Phi_{i}\right),$
$L_{cos}(x,y)$ $=1-\frac{x^{\intercal}y}{\left|x\right|\left|y\right|}$ (2)
$L_{summary}(x)$ $=\sum_{i}\lambda_{i}L_{cos}(y_{i}^{(s)},z_{i}^{(s)})$
We found empirically that cosine distance loss produced better models compared to L1, MSE, Smooth-L1 [19].
We found that additionally supervising the spatial features of the model by matching the teacher was not only important for downstream dense tasks, it also improved the holistic quality of our model.
Teachers Zero Shot k-NN ADE20K
None 75.77 82.59 41.18
CLIP 75.64 82.60 44.42
DINO 74.68 83.02 47.05
Both 74.85 82.96 48.13
Table 2: Ablation over which teachers we supervise the spatial features. We use a ViT-L/14 student model and train on the LAION-400M dataset. Adding this loss term is always beneficial. Dino appears to provide better spatial features than CLIP, but training the student to match both teachers produces the best results.
For matching the spatial features, we employ a combination of cosine similarity and smooth L1. Similar to equation (2) where we found that cosine similarity produced the best results, we found the same to be true for the spatial features, however, since we want to allow our student model to be a drop-in replacement in the teacher frameworks, it’s also important that we match the magnitude of the teacher vectors, and so we introduce smooth L1. In (3) we show the formulation of this loss. Let $h_{i}^{(v)}(x^{\prime}|\Theta_{i}^{(v)})$ be the learned student head for matching teacher feature vectors, and corresponding $t_{i}^{(v)}(x|\Phi_{i}^{(v}))$ be the teacher feature vectors, with $x^{\prime}=f(x|\Theta_{0})$, then the spatial feature loss is:
$L_{\text{match}}(x,y)$ $=\alpha L_{cos}(x,y)+\beta L_{smooth_{l}1}(x,y)$ (3)
$L_{\text{features}}(x)$ $=\sum_{i}\gamma_{i}L_{match}\left(h_{i}^{(v)}(x^{\prime}|\Theta_{i}^{(v)}),t_{i}^{(v)}(x|\Phi_{i}^{(v}))\right)$
We choose $\alpha=0.9$ and $\beta=0.1$ to mostly rely on the empirically better cosine distance, but to also match vector magnitudes.
3.3.1 Loss Balancing
Due to the number of possible combinations of loss weights between the different teachers, and even which teachers, and possible formulations of loss functions, we mostly opted toward naive loss balancing with all teachers equally weighted for spatial features. For summary features, we have $\lambda_{CLIP}=\lambda_{DINO}=1$ and $\lambda_{SAM}=0.1$ owing largely to SAM’s low performance on k-NN classification. We then also set $\gamma_{i}=1$ by default.
We did experiment with automatic loss balancing using predicted uncertainty [10], AdaLoss [24] (momentum 0.99) and separately with AMTML-KD [37], as ways to learn the balance of $\lambda_{i}$ and $\gamma_{i}$. In the case of AMTML-KD, the model would always collapse its entire weight around the CLIP teacher and would yield worse results than naive manual balancing. Based on the results in table 3, there is very little advantage to the more exotic balancing schemes, so we opt for the ”Naive” method throughout the rest of the paper.
Method Zero Shot k-NN ADE20K
Naive 70.63 79.50 44.71
Uncertainty [10] 70.92 79.37 44.57
AdaLoss [24] 71.31 79.77 44.36
Table 3: Loss term balancing methods comparison. We use a ViT-B/14 student, and CLIP+DINOv2 teachers. We found that AdaLoss produces the best results on the ImageNet tasks, but the worst on ADE20K. Ultimately, they’re all really close.
4 Implementation Details
Performing multi-teacher distillation is not trivial due to a mismatch in feature dimensions, different input resolution, different concepts for loss computation, challenges in fitting multiple teachers into a single GPU. We address these issues next.
General. We train all student models using the AdamW [40] optimizer, batch size 1024, and cosine annealing learning rate schedule and base learning rate of $31e-$ 3. We train for 600k steps, resulting in 614M total examples seen, approximately 2% of the number of examples seen by OpenCLIP models for training from scratch. We get our CLIP teacher models from the OpenCLIP model zoo [^1] https://github.com/mlfoundations/open_clip, our DINO-V2 models from their model zoo  [^2] https://github.com/facebookresearch/dinov2, and SAM from their model zoo  [^3] https://github.com/facebookresearch/segment-anything. For our best student models, we train using the MetaCLIP [59] ViT-H/14 OpenCLIP model, the ViT-g/14 DINO-V2 model, and the Huge SAM model. We apply random scale + cropping  [^4] https://timm.fast.ai/random_resized_crop to both student and teacher inputs (e.g. all models see the same image). We chose the DataComp-1B dataset due to it having the highest quality results of the web-scale datasets we had access to.
Student architecture. We study two settings for student model architecture:
  • Standard ViT [14] architecture to match the architecture of teachers.
  • Efficient architecture variants prioritize high throughput on GPUs. See Section 5.1.
Multi-scale Teachers. Both CLIP and DINO primarily operate at a resolution of $224^{2}$, which makes it trivial to employ both to train a student model at the same resolution. SAM, however, operates at an input resolution of $1024^{2}$, which forces us to employ a few techniques.
Rank/Teacher Partitioning. We group teacher models by (batch_size, resolution), and then distribute the groups to different distributed ranks, such that each GPU processes a consistent batch size and input resolution. We also sample groups at different rates. For our training setups that include SAM, we train with 64 GPUs (ranks), half of which get the CLIP+DINO group with batch size 32 per GPU and input resolution 224, and the other half get SAM with batch size 4 per GPU and input resolution 1024. This results in an effective batch size of 1,152. For CLIP+DINO training, we use 32 GPUs, resulting in batch size 1024.
Multi-Resolution ViTs. Many of our student models use ViT [14] as the base vision architecture. Traditionally, ViTs use a learned position embedding for each input patch in an image, which in turn enforces that the model always operate at a constant resolution. We employ the Cropped Position Embedding (CPE) [29] augmentation with the number of positions being equal to $64^{2}$ or the size of the teacher, whichever is larger. The position embeddings are then randomly cropped and interpolated to match the number of input patches for the student model. Even when training with CLIP+DINO at 224 resolution, we found that this technique results in a negligible drop (Table 4) in holistic metrics, but improved semantic segmentation linear probing mIOU. For heterogeneous-resolution teachers, this is a seamless technique that allows ViT to operate at arbitrary resolutions within some envelope.
Method k-NN ADE20K
OpenCLIP H/14 81.10 40.04
DINO-v2 L/14 83.44 46.55
RADIO ViT-L/14 Non-CPE 82.96 47.30
RADIO ViT-L/14 CPE 82.84 48.52
Table 4: Comparing identical RADIO-ViT models, with and without the CPE [29] formulation, distilled from DINO-v2 ViT L/14 and OpenCLIP ViT H/14 teachers. For ADE20K segmentation, we use a simple linear head setup with the image processed at $518^{2}$ resolution. Our models were trained using the LAION-400M dataset for this experiment. Notably, the student model only ever trains at $224^{2}$ resolution, but is able to generalize to $518^{2}$ and even outperform its DINO-v2 teacher on semantic segmentation.
High-Resolution ViT Student. In SAM, they employ the ViTDet [33] architecture as a way to reduce the computational and memory burden of ViT models at high-resolution. A key part of the architecture is to convert the MHSA layers to windowed attention, and then have a normal global attention layer every $k$ layers. It turns out that this architecture choice can easily be adapted into a training augmentation, where we sample a window size from a set of possible window sizes, and for a given step, the windowed layers have that size. This allows us to meaningfully reduce the computational burden of training the student model with the SAM teacher, and, as we make the window size flexible, it provides an additional scaling mechanism during inference where the same model may be used at different degrees of computation. Table 8 shows the resulting quality of the same model being run at different window sizes.
Student/Teacher Downsampling Mismatch. When the student and teacher downsample images through their processing stack at different rates, it results in the output feature vectors having different resolutions. For example, if the teachers use a ViT-H/14 architecture and student a ViT-H/16, it means that the student outputs a $14^{2}$ feature map, and the teachers a $16^{2}$ feature map. For $L_{\text{features}}$ we bilinearly interpolate the student outputs to match the resolution of the teacher features in order to apply the loss to each output vector.
Summary Loss Feature Zero Shot k-NN ADE20K VOC GQA SQA TextVQA VQAv2
CLS token 0.07 0.09 78.55 83.91 49.01 83.51 70.46 65.89 53.36 77.66
Avgpool 0.05 0.12 80.12 83.83 38.36 77.04 70.92 65.84 54.26 78.28
Table 5: Comparing identical ViT models, with CLS token and average pooling summarization. Summary Loss - training summary loss against CLIP teacher, Feature - training feature loss against DINOv2 teacher.
Feature Summarization. In 3.3 we explained how teacher summary features are extracted using the “class token” of their respective ViT models. We now turn our attention to the summarization of student features. ViTs have 2 options: (i) a separate summarization “CLS” token or (ii) average pooling patch tokens. We evaluate both options in Table 5. We observe that average pooling improves summary loss, but has a more significant detrimental effect on the feature loss. Given the importance of the latter we choose to use separate CLS tokens in the rest of this study.
5 Results
In this section, we analyze models obtained with the proposed AM-RADIO framework. First, we touch upon backbone efficiency, then compare with the original teachers (CLIP, DINOv2, SAM), and benchmark models under vision question answering in the LLaVa framework. We will see that the proposed models outperform the original teachers in multiple metrics, including throughput. A sneak peek of the results is presented in Figure 1.
Refer to caption Refer to caption Refer to caption
Benchmarking multiple model architectures in AM-RADIO pipeline ADE20K segmentation example
Figure 2: All models followed the same training protocol. The results from three benchmarks show that RADIO and E-RADIO models outperform others in efficiency. This underperformance in other models might be due to overfitting architectures on supervised ImageNet-1K training. E-RADIO notably delivers results 10 times faster and with a 20% improvement over teacher models. Segmentation results demonstrate that SAM lags behind DINOv2 in segmentation tasks, likely because it is tasked with learning multiple plausible masks, which requires features for object interiors to vary relative to edges.
5.1 Efficient Students
Backbone Param. Through- Zero k-NN ADE20k FD loss
Count put Shot
Teachers
DINOv2 G/14 1.14B 313 N/A 83.41 47.53
OpenCLIP H/14 632M 556 77.19 81.10 40.04
ViT models
RADIO-ViT B/32 88M 17152 70.39 75.64 36.15 0.382
RADIO-ViT B/14 86M 3699 73.84 80.73 44.71 0.190
RADIO-ViT L/14* 304M 1126 77.25 84.03 48.7 0.151
RADIO-ViT H/14* 632M 556 78.62 84.17 49.01 0.087
Existing efficient models
EfficientNetV2-S 21M 9017 65.37 70.72 27.75 0.415
ResNetv2-101 44M 7283 69.58 75.32 29.61 0.405
RegNetY-064 30M 6573 69.84 74.59 28.9 0.394
EfficientViT-L1 38M 6048 71.73 79.90 33.12 0.376
ConvNext-B 88M 1805 75.43 81.73 38.95 0.358
NFNet-F3 254M 1777 76.93 80.50 38.31 0.340
SwinV2-S 49M 1497 74.70 81.12 35.57 0.364
MaxViT-B 119M 1486 77.49 79.34 38.46 0.340
PoolformerV2-M36 56M 1194 74.46 80.49 35.05 0.377
MViTV2-B 51M 975 75.92 81.39 41.39 0.345
Proposed architecture
E-RADIO-B 118M 6422 75.19 82.21 44.03 0.319
[…]
E-RADIO-L 265M 3472 77.87 83.73 45.5 0.265
Table 6: Comparison of backbones. Throughput is measured using TensorRT 9.0.1 on A100 in mixed FP16/FP32 precision at batch size 128 on 224x224 resolution. Rows are sorted by descending throughput order. FD loss refers to the Feature Distillation training loss against the DINOv2 teacher, which exhibits high correlation with the ADE20k mIoU. Bolded models are on the speed/quality Pareto front.
We aim to find an efficient model architecture to speed up the inference of VFM. There are a number of architectural designs aimed at high throughput on GPU devices. Unfortunately, most studies employ varying training recipes, even when training and evaluating on the ImageNet-1K dataset, making comparisons misleading. We use our distillation framework to evaluate several backbones with no change in training hyperparameters.
Upon reviewing the literature on efficient vision backbones focused for high GPU throughput, we pick the following list of architectures: EfficientNetV2[53], ResNetv2[52], RegNetY[47], FasterViT[21], EfficientViT[7], ConvNext[39], NFNet[6], SwinV2[38], MaxViT[54], PoolformerV2[63] and MViTV2[34]. We train all the backbones via distillation on the ImageNet-21k dataset, using OpenCLIP ViT-H/14 (laion2B-s32B-b79K) and DINOv2 G/14 as teachers. We train for 600k steps at a total batch size of 4096, which is empirically deemed enough to reach convergence. Results are compiled in Table 6.
We observe that many models lag behind teachers. Additionally, CNN-like models are significantly faster than ViTs, while the latter are more accurate. The relatively low performance of existing efficient backbones on the dense ADE20k segmentation task is not unexpected since all of them apply a spatial dimension reduction factor of 32 for a final feature maps: the size is 7x7px for input resolution of 224x224px. Thus hardly capable of capturing fine-grain spatial information.
E-RADIO: To overcome this issue, we propose a novel hybrid architecture, named E-RADIO (Efficient RADIO). This design borrows ideas from existing literature and includes an input stem with strided convolutions to downsample the input image by 4x. It then proceeds with 2 stages of YOLOv8 C2f convolution blocks and 2 stages of transformer. For the transformer variant we pick windowed attention (like in SWIN [38]), and interleave local windowed attention with “global” windowed attention as done in [21] and ViTDet [33]. To perform “global” attention we first downsample the feature map by 2x, apply windowed attention, and then upsample the feature maps back to the original resolution. Up-/down-sampling is performed by strided convolution with a kernel size 3x3 and a stride of 2. The last idea is borrowed from EdgeViT [44], which uses local-global-local attention. Details of the architecture are in the Appendix.
Feature map upsample: E-RADIO upsamples final features maps by 2x via a deconvolutional layer and adds them to feature maps from the third stage, resulting in only a 16x spatial resolution reduction. Such upsampling gives an improvement in dense task while being only 10% slower.
Results of E-RADIO in Table 6 demonstrate that the proposed architecture significantly outperforms the competition, and can be seen as an efficient replacement for the much slower full ViT. We pick E-RADIO as our efficient candidate for further study. We envision our framework to be a benchmarking suite for future efficient vision backbone research.
Family Arch Throughput Zero-Shot k-NN ADE20k VOC GQA SQA TextVQA VQAv2
OpenCLIP [9] ViT-H/14 556 77.19 81.10 40.04 68.03 57.86 67.53 48.11 67.31
MetaCLIP [59] ViT-H/14 556 80.51 82.12 35.39 62.62 67.85 68.77 52.02 75.46
OpenAI CLIP [46] ViT-L/14 1126 75.54 79.80 36.51 67.04 68.80 67.03 53.25 76.47
DINOv2 [43] ViT-L/14 1126 N/A 83.44 46.55 82.07 70.16 63.76 45.49 75.37
ViT-g/14 344 N/A 83.41 47.53 82.95 69.29 64.60 45.25 74.56
SAM [31] ViTDet-L/16 14 N/A 20.41 26.56 34.34 61.52 63.36 43.29 64.91
RADIO (Ours) E-RADIO-L 3697 77.87 83.73 45.50 79.95 66.22 65.54 43.51 70.74
${}^{\dagger}$ 556 82.73 85.29 50.32 85.17 69.17 67.08 54.92 77.88
Table 7: Comparison of teachers and RADIO models. ”Zero-Shot” and k-NN are computed on ImageNet-1K. ADE20K and VOC refer to linear probe semantic segmentation mIOU on the ADE20K [68] and PascalVOC2012 validation datasets. RADIO outperforms all teacher models on most metrics, and then E-RADIO-L is about 6.6x faster than ViT-H/14 while attaining similar metrics on certain tasks.
${}^{\dagger}$e released RADIOv1 model weights after initial publication. We used DFN CLIP [16] (84.37 Zero-Shot), OpenAI CLIP [46], and DINOv2 [43] to train this model. It additionally incorporates 5 register tokens [12]. Refer to table 11 for the original publication numbers.
5.2 Comparison with teachers
A comprehensive set of results is presented in Table 7. At first, we evaluate teacher models: OpenCLIP, MetaCLIP, and DINOv2. We notice that MetaCLIP is better than OpenCLIP and therefore choose it. DINOv2 provides important properties for dense tasks: ADE20k and VOC. Our E-RADIO-L model is 6.6x faster than MetaCLIP and 11.1x faster than DINOv2-ViT-g/14. At the same time, it outperforms MetaCLIP on k-NN, ADE20k, VOC, and GQA, while also enabling Zero-shot capability that is absent in DINOv2. Our full model, RADIO-ViT-H/14, is as fast as the teachers but outperforms them on 6 out of 8 tasks, demonstrating the efficiency of the proposed distillation framework. Details of these experiments can be found in the next sections.
5.3 Drop-In SAM Replacement
Following [7], we use their evaluation harness to compute the mIOU for instance segmentation using pretrained SAM with vision encoder replaced by our model. Table 8 shows the results of the COCO Instance Segmentation task using the baseline SAM models, and RADIO-ViT.
COCO 2017 drop-in SAM replacement at 1024x1024
Family Arch mIOU Throughput
SAM Base 75.78 29.18
Large 77.01 13.52
Huge 77.18 8.56
RADIO (ours) ${}^{\dagger}$ 75.17 8.20
${}^{\dagger}$ 75.31 8.04
${}^{\dagger}$ 75.42 7.45
Other benchmarks
Family Arch Zero-Shot k-NN ADE20k
SAM ViTDet-L/16 20.41 26.56 34.34
ViTDet-H/16 22.39 28.17 32.16
RADIO (ours) ViTDet-H/14 76.17 83.34 49.01
Table 8: In our SAM replacement experiments using the COCO 2017 instance segmentation framework, we substitute SAM’s vision encoder with our agglomerative (RADIO) model. This model aligns with features just before the encoder’s bottleneck layer. We also examine the impact of varying ViTDet window sizes. Our model almost matches SAM’s results, primarily as a multi-teacher distillation byproduct. Further, when comparing metrics on ImageNet1K and ADE20k, our RADIO model shows superiority.
5.4 Comparison with SAM-CLIP [55]
Concurrently with our work, SAM-CLIP was introduced as a method of fusing SAM and CLIP into a single model. Due to the concurrency of effort, we don’t compare our model with the full suite of metrics demonstrated in their method, however, we do have some overlap in key metrics such as Zero-Shot ImageNet1k, and ADE20k semantic segmentation via linear probing. We present the comparison in table 9, however we note that there are enough differences between these two models that we can’t conclude one way or another what is the superior approach. Instead we’ll argue that DINOv2 does a better job of ADE20k linear probing than SAM, and thus our significantly higher quality on this metric is likely due to the inclusion of DINOv2, which is a key introduction with our approach.
Family Model Zero-Shot ADE20k
SAM ViTDet-H/16 28.2
CLIP ViT-L/14 79.2 36.4
SAM-CLIP ViTDet-B/16 71.7 38.4
RADIO ViT-H/14 78.6 49.0
Table 9: We compare our common key metrics with those demonstrated in SAM-CLIP [55]. We note that there are numerous differences between the two approaches, including model capacity and architecture. SAM-CLIP uses the ViT-B variant of SAM as a starting point, which implies it’s a ViTDet-B/16 architecture. As a result of this choice, their metrics are computed at a resolution of 1024x1024. RADIO trains a vanilla ViT-H/14 from scratch, and as a result of the flexibility gained via the CPE method, we evaluate Zero-Shot ImageNet1k at a resolution of 224x224, and we run ADE20k linear probing at a resolution of 518x518 using the exact same weights. We note that Zero-Shot quality is largely determined by the quality of the CLIP teacher and the capacity of the student. We attribute our superior quality on ADE20k semantic segmentation largely to our inclusion of DINOv2 as a teacher.
Figure 3: VQA examples from LLaVa-1.5 framework with swapped vision encoder. RADIO and E-RADIO are better at capturing fine details, and it is better at counting. We observed that MetaCLIP, DINO and RADIO have high correlation of answers at the level of 0.78, while E-RADIO has 0.68 with others and tends to choose a different answer.
5.5 Semantic Segmentation Linear Probing
We train a linear head on top of the frozen features of the teachers and students alike and evaluate performance in the MMSeg [11] framework using the mIoU metric on ADE20k and PascalVOC2012 datasets. We use a training and evaluation crop size of 518x518 for models that support this resolution, or 224x224 for others. We use the “slide” evaluation mode with a stride of $\frac{2}{3}$ the crop size. We train the linear head for 160k steps using a total batch size of 16, a base learning rate of $10^{-3}$ and the AdamW optimizer.
5.6 Visual Question Answering
We replace the vision encoder in a LLaVA 1.5[35] setup with our own encoder. A 2-layer MLP is used to project frozen visual features into the language token space. Under the default LLaVA 1.5 settings, we pretrain a multimodal projection MLP and then run instruction tuning to finetune a Vicuna 7B-1.5 model[67]. We evaluate models using the validation sets of GQA [26], TextVQA [50] and ScienceQA [41] (image subset), and we score the model on the Test-Dev set of VQAv2 [20] using EvalAI[60]. We use an input image resolution of 336x336 when the vision encoder model supports it, and fall back to 224x224 otherwise. Experimental results are compiled in Table 7. Qualitative results are shown in Figure 3.
6 Key Insights
Outperforming teachers. Most Vision Foundation Models (VFMs) have unique properties such as language grounding (CLIP), dense correspondences (DINOv2), and detailed segmentation (SAM). Distillation allows uniting all these properties in a single model that even outperforms each of them. We have also observed that better teachers yield better students, which allows RADIO to absorb the current SOTA foundation models at a given point in time.
Training efficiency. Our distillation framework consumes only 2%-5% of the data samples compared to training CLIP [18] or DINO from scratch. Even with the costs of running multiple teachers during training, the overall compute is only a quarter of that required for training CLIP from scratch.
Feature distillation loss. We observe the crucial importance of full feature distillation to boost the performance of the teacher in dense image understanding tasks, such as an 18% relative improvement on ADE20K. Future research focusing on developing new loss formulations could further enhance performance.
SAM vs DINOv2. We find that, out of the box, SAM is not well-suited for downstream tasks, whereas DINOv2 significantly outperforms in zero- and few-shot tasks. For example, ADE20K segmentation via linear probing is 1.7x better with the latter, and the ImageNet1k k-NN metric is 4x better. SAM excels in detecting edges and segmenting objects but performs poorly in high-level object description and combining the semantics of multiple objects (Figure 2).
Efficient backbone. Most of the existing VFMs use ViTs as the model architecture for the vision encoder. Based on our analysis of distilling efficient backbones, we conclude that most model designs are overly tailored towards supervised training on ImageNet1K and, as a result, do not scale well to VFM settings due to the low feature resolution at the output. We designed a new vision backbone, E-RADIO, with a hybrid CNN-Transformer architecture that significantly outperforms the competition. While it is at least 6x faster, it still lags behind pure transformers on dense tasks.
Conclusions
We propose a multi-teacher distillation framework to efficiently train vision foundation models (VFMs). We demonstrate that it can unify unique properties of other VFMs such as CLIP, DINOv2, and SAM. Our models outperform the original teachers in multiple tasks under the same model architectures while exhibiting greater capability. On the efficiency side, we designed an architecture, E-RADIO, that is up to 10x faster, while closely matching or even outperforming the teachers. We believe our work will advance the field of efficient VFM design and benchmarking.
Acknowledgment
We would like to thank Hongxu (Danny) Yin, Karan Sapra, and Francesco Ferroni for their important feedback on the project and Philipp Fischer and Lukas Vögtle for help with training infrastructure.
Supplementary Material
Appendix A Teacher Comparison
Family Arch Throughput Zero-Shot k-NN ADE20k VOC GQA SQA TextVQA VQAv2
OpenCLIP [9] ViT-H/14 556 77.19 81.10 40.04 68.03 57.86 67.53 48.11
MetaCLIP [59] ViT-H/14 556 80.51 82.12 35.39 62.62 67.85 68.77 52.02 75.46
DINOv2 [43] ViT-L/14 1126 N/A 83.44 46.55 82.07 70.16 63.76 45.49 75.37
ViT-G/14 344 N/A 83.41 47.53 82.95 69.29 64.60 45.25 74.56
SAM [31] ViTDet-L/16 26 N/A 20.41 26.56 34.34 59.87 67.08 43.90 62.46
ViTDet-H/16 16 N/A 22.39 28.17 32.16 57.62 63.26 43.34 60.00
Table 10: Comparison of teachers models. ”Zero-Shot” and k-NN are computed on ImageNet-1K. ADE20K and VOC refer to linear probe semantic segmentation mIoU on the ADE20K [68] and PascalVOC2012 validation datasets. Note: for SAM, throughput and all other metrics were measured using the only supported image resolution of 1024x1024. We report k-NN metrics and Visual Question Answering average pooling the 64x64 patch tokens into a 16x16 feature map, and emphasize that SAM was not designed for these tasks. For other models, we used an image resolution of 224x224 for image classification, 336x336 for visual QA and 518x518 for semantic segmentation.
Figure 4: Comparison of teachers within our suite of downstream evaluation tasks.
Teacher metrics are summarized in Table 10. Interestingly, we note that even though MetaCLIP performs better than OpenCLIP on image-level classification and visual question answering metrics, it exhibits lower performance on semantic segmentation.
Appendix B E-RADIO architecture details
Figure 5: High level architecture of the ERADIO network architecture. Overall architecture is composed of multiple stages: 1) the stem, 2) 2 convolutional blocks from YOLOv8, 3) 2 transformer blocks with multi-resolution windowed self attention.
The architecture of E-RADIO is illustrated in Figure 5. It is a hybrid CNN-Transformer architecture. First 2 stages follow convolution paradigm and have the C2f architecture from YOLOv8 model [28]. The last 2 stages have the Transformer architecture with windowed attention and multi-resolution attention (MRA) structure. Every stage, except the last one, are followed by downsample block. We implement it as a strided convolution with 3x3 kernel and stride 2, followed by batch normalization layer.
B.1 Multi-Resolution Attention
Figure 6: Multi-resolution attention for E-RADIO
Standard transformers struggle to scale with high input image resolution because of quadratic complexity of the attention. SWIN [38] proposed to use windowed attention to reduce the complexity of attention. We reuse windowed attention in the E-RADIO. To address for missing communication between windows, SWIN introduced window shifting, unfortunately, it has non-negligible compute cost. Instead, we propose multi-resolution attention inspired by EdgeViT’s Local-Global-Local attention [44]. The idea is illustrated in Figure 6. Every layer in the transformer will have a local windowed attention with optional subsampling via convolutional operator. For example, if susbampling is dissabled, then it is just a standard windowed attention. If the subsampling ratio is 2, then the feature map is downsampled by a factor of 2, windowed attention is performed, and then the feature map is upsampled to the original resolution with deconvolution. For FasterVIT2 models, we interleave subsampled attention with ratio 2 and the normal attention with no subsampling.
B.2 Configurations
All models in the family follow the same configuration except the embedding dimension (hide dimension). We simply scale it up with bigger models. Other parameters:
  • Input resolution is 224
  • In-stem contains 2 3x3 convolutions with stride 2
  • Total stages: 2 convolutional and 2 transformer
  • First stage takes input feature size of 56x56, has 3 layers with C2f structure from YOLO8 [28].
  • Second stage takes input feature size of 28x28, has 3 layers of C2f.
  • Third stage takes features of size 14x14, has 5x multi-resolution attention, window size 7.
  • Forth stage takes features of size 7x7, has 5x windowed attention of window size 7.
  • Embedding dimension for different model variants: XT - 64, T - 80, S - 96, B - 128, L - 192. The smallest XT and T models have [1, 3, 4, 5] layers for each of 4 stages.
  • Output features have resolution of 14x14 and are obtained by upsampling the features of stage 4 by 2x with deconvolution and adding to stage 3 features of size 14x14.
Appendix C ViTDet Augmentation
The following python code shows how the alternating window/global architecture of ViTDet [33] can be applied to a transformer. We take advantage of the fact that transformers are permutation invariant after position encodings have been applied, and thus it’s easy to organize the patch order such that contiguous chunks of patches belong to the same window. Once reordered in this way, alternating between windowed and global attention is achieved simply by absorbing the windows into the batch dimension or returning to the original shape respectively. We also enforce that the final transformer layer always applies global attention.
from einops import rearrange
def reorder_patches(patches: torch.Tensor,
                    patched_size: Tuple[int, int],
                    window_size: int):
    p_idxs = torch.arange(patches.shape[1])
    p_idxs = rearrange(p_idxs, ’(wy y wx x) -> (wy wx y x)’,
                       wy=patched_size[0] // window_size, y=window_size,
                       wx=patched_size[1] // window_size, x=window_size)
    p_idxs = p_idxs.reshape(1, -1, 1).expand_as(patches)
    return torch.gather(patches, p_idxs), p_idxs
def vitdet_aug(blocks: nn.Sequential,
               patches: torch.Tensor,
               patched_size: Tuple[int, int],
               window_sizes: List[int],
               num_windowed: int):
    B, T, C = patches.shape
    window_size = sample(window_sizes)
    sq_window_size = window_size ** 2
    patches, p_idxs = reorder_patches(patches, patched_size, window_size)
    period = num_windowed + 1
    for i, block in enumerate(blocks[:-1]):
        if i % period == 0:
            patches = patches.reshape(B * sq_window_size, -1, C)
        elif i % period == num_windowed:
            patches = patches.reshape(B, T, C)
        patches = block(patches)
    # Always use global attention with the last block
    patches = patches.reshape(B, T, C)
    patches = blocks[-1](patches)
    # Finally, put the patches back in input order
    ret = torch.empty_like(patches)
    ret = ret.scatter(dim=1, index=p_idxs, src=patches)
    return ret
Appendix D Automatic Loss Balancing
D.1 Uncertainty
Following [10], we have:
$L(x)=\sum_{k}\frac{1}{2\sigma_{k}^{2}}L_{k}(x)+\log\sigma_{k}$ (4)
where the $\sigma_{k}$ values are predicted by the student. In practice, the student predicts $b:=\log\sigma_{k}^{2}$ for numerical stability, to avoid division by zero, and to regress unconstrained scalar values.
We make some minor modifications to (4) to make training a bit more stable in our setting. We replace the manual $\lambda$ scalars with the learned uncertainty weights, and add the loss term for large uncertainties. Altogether, this yields:
$\lambda_{k}$ $=\frac{e^{-b_{k}}}{2}$ (5)
$L(x)$ $=\sum_{k}\lambda_{k}L_{k}(x)+\frac{b_{k}}{2}$
Let $b_{i}^{(s|v)}(x^{\prime}|\Theta_{i}^{(s)})$ be a learned function predicting balance parameters for teacher $i$ and summary weight $(s)$ or feature vector weight $(v)$, we transform equation (5) slighty to:
$\psi(x)$ $=\log(1+e^{x})$ (6)
$\lambda_{i}^{(m)}$ $=e^{-b_{i}^{(m)}(x^{\prime})}$
$L(x)$ $=\sum_{i}\sum_{m\in\{s,v\}}\lambda_{i}^{(m)}L_{i}^{(m)}(x)+\psi\left(b_{i}^{(m)}(x^{\prime})\right)$
The function $\psi(x)$ is the familiar ”softplus” nonlinear activation function. We drop the division by 2 on the left because, assuming outputs are initially $b\sim\mathcal{N}(0,\sigma^{2})$, then the loss weights will initially have an expected value of 1, matching the naive weighting. On the right, we replace $\frac{b_{k}}{2}$ with $\psi(x)$ for a few reasons:
  • When $x\gtrapprox 4$, then $\psi(x)\approx x$, yielding the same expression as before.
  • When $x\approx 0$, then $\psi^{\prime}(x)\approx\frac{1}{2}$, yielding the same expression as before.
  • When $x<0$, which translates to a loss weight $>1$, $\psi^{\prime}(x)\to 0$, improving stability as the weight gets larger.
  • It has range $(0,\infty)$ which aesthetically enforces the loss to be greater than zero.
D.2 AdaLoss
In addition to uncertainty auto-balancing, we also explored AdaLoss [24]. In this formulation, we have:
$\lambda_{i}^{(m)}$ $=\frac{1}{\mathbb{E}(L_{i}^{(m)})}$ (7)
$L(x)$ $=\sum_{i}\sum_{m\in\{s,v\}}\lambda_{i}^{(m)}L_{i}^{(m)}(x)$
Appendix E Visual Question Answering Samples
Figure 7: Visualization of the LLaVA attention maps over the visual features produced by a RADIO encoder. We use one sample image from the GQA[26] validation set and one associated question: ”What color is the helmet in the middle of the image?”. For each layer in the language model, we retrieve attention scores for all positions of the visual tokens, average them over all attention heads, and overlay corresponding heat maps with the input image. We can see that as we progress through the layers, the model’s attention focuses on the relevant part of the image. The model’s answer is ”Blue”.
Figure 8: Sample questions from the GQA[26] and their answers from our LLaVA models, using various image encoders. Answers are painted green when they match the ground truth, pink otherwise.
Figure 9: Sample questions from the GQA[26] and their answers from our LLaVA models, using various image encoders. Answers are painted green when they match the ground truth, pink otherwise.
Figure 10: Sample questions from the GQA[26] and their answers from our LLaVA models, using various image encoders. Answers are painted green when they match the ground truth, pink otherwise.
Figure 11: Sample questions from the TextVQA [50] dataset and their answers from our LLaVA models, using various image encoders. Answers are painted green when they match the ground truth, pink otherwise.
Figure 12: Sample questions from the TextVQA [50] dataset and their answers from our LLaVA models, using various image encoders. Answers are painted green when they match the ground truth, pink otherwise.
Figure 13: Sample questions from the TextVQA [50] dataset and their answers from our LLaVA models, using various image encoders. Answers are painted green when they match the ground truth, pink otherwise.
Figures 8 to 13 show sample questions from our Visual Question Answering datasets, together with sample answers when using our vision encoders in a LLaVA setup.
Appendix F Original Model Comparison Table
Table 7 shows updated RADIO results after the original publication date. For reference, we provide the original table as table 11.
Family Arch Throughput Zero-Shot k-NN ADE20k VOC GQA SQA TextVQA VQAv2
OpenCLIP [9] ViT-H/14 556 77.19 81.10 40.04 68.03 57.86 67.53 48.11 67.31
MetaCLIP [59] ViT-H/14 556 80.51 82.12 35.39 62.62 67.85 68.77 52.02 75.46
OpenAI CLIP [46] ViT-L/14 1126 75.54 79.80 36.51 67.04 68.80 67.03 53.25 76.47
DINOv2 [43] ViT-L/14 1126 N/A 83.44 46.55 82.07 70.16 63.76 45.49 75.37
ViT-g/14 344 N/A 83.41 47.53 82.95 69.29 64.60 45.25 74.56
SAM [31] ViTDet-L/16 14 N/A 20.41 26.56 34.34 61.52 63.36 43.29 64.91
RADIO (Ours) E-RADIO-L 3697 77.87 83.73 45.50 79.95 66.22 65.54 43.51 70.74
ViT-H/14 556 78.62 84.17 49.01 83.51 70.46 65.89 53.36 77.66
Table 11: Comparison of teachers and RADIO models. ”Zero-Shot” and k-NN are computed on ImageNet-1K. ADE20K and VOC refer to linear probe semantic segmentation mIOU on the ADE20K [68] and PascalVOC2012 validation datasets. RADIO outperforms all teacher models on most metrics, and then E-RADIO-L is about 6.6x faster than ViT-H/14 while attaining similar metrics on certain tasks.